The angles of a triangle are (2x - 30)degree,(2x -60)degree and (x-30)degree.what is the value of x?
Answer by Cromlix(4381)

Total degrees in a triangle = 180 degrees
2x - 30 + 2x - 60 + x - 30 = 180
Collect like terms
5x - 120 = 180
5x = 180 + 120
5x = 300
x = 60 degrees.
Hope this helps.
:-)